OpenStudy (anonymous):

Prove that the magnitude of the resultant of two equal forces in magnitude acting at one point and the angle between them is 20 , is equal in magnitude to each force

Parth (parthkohli):

Do you mean 120?

Parth (parthkohli):

\[|C| = \sqrt{|A|^2 + |B|^2 + 2|A||B|\cos(\theta)}\]Well, that formula's a start.

OpenStudy (anonymous):

Yes Im sorry

OpenStudy (anonymous):

Then ?

Parth (parthkohli):

Oh, sorry. Uh -- so the magnitudes of \(A\) and \(B\) are equal and \(\theta = 120\).

OpenStudy (anonymous):

Okaai I did that And then how can i get the two resultant ?

Parth (parthkohli):

Use the value \(\cos(120) = \dfrac{-1}{2}\)

OpenStudy (anonymous):

i did that :)

OpenStudy (anonymous):

Please write the complete answer cause i want to ask alot of questions

Parth (parthkohli):

\[|C| = \sqrt{A^2 + A^2 + 2|A|^2\cdot \dfrac{-1}{2}} = \sqrt{|A|^2}\]

OpenStudy (anonymous):

hmm

Parth (parthkohli):

Uh, wouldn't that... suffice?

Parth (parthkohli):

What is \(\sqrt{|A|^2}\)

OpenStudy (anonymous):

I dont know :/

Parth (parthkohli):

What is the square root of the square?

OpenStudy (anonymous):

i dont understand you